# Stochastic theta methods for random periodic solution of stochastic differential equations under non-globally Lipschitz conditions11footnotemark: 1

### v1

@@ 5 Numerical experiments
$~{}+a\pi(|x||y|^{3}+|y|^{4})|t-s|+6\pi^{2}d^{2}(p^{*}-1)|y|^{4}|t-s|$
$\leq$ $~{}3c^{2}(p^{*}-1)|x-y|^{2}+\big{(}2a\pi+6\pi^{2}d^{2}(p^{*}-1)\big{)}(1+|x|^{4}+|y|^{4})|t-s|,$
which together with (5) indicates that (A3) holds with $L_{f}=3c^{2}(p^{*}-1)$ and $\gamma=3$. In combination with (5.4), $L_{f}\in(0,\lambda)$ and $p^{*}>4\gamma-2$, we choose
@@
One can check that the associated period is $1$ and Assumptions 2.1 and 4.3 are fulfilled. It is worth noting that (5.5) has been considered in [22] to test the periodicity and mean square convergence of the backward Euler method. Thus for simplicity, we only verify the mean square order of ST methods on the time interval $[-10,10]$ with $\theta=0.75,1$ under the above settings for convergence test of (5.1). Figure 5 shows that the slopes of the error lines and the reference lines match well, indicating that the ST methods have strong rates of order one order one in additive case. Actually, least square fits produce rate $1.0289$ for $\theta=0.75$ and rate $1.0289$ for $\theta=1$. These facts coincide with the previous theoretical result in Theorem 4.6.

### v2

@@ 5 Numerical experiments
$~{}+a\pi(|x||y|^{3}+|y|^{4})|t-s|+6{\pi}d^{2}(p^{*}-1)|y|^{4}|t-s|$
$\leq$ $~{}3c^{2}(p^{*}-1)|x-y|^{2}+\big{(}2a\pi+6{\pi}d^{2}(p^{*}-1)\big{)}(1+|x|^{4}+|y|^{4})|t-s|,$
which together with (5) indicates that (A3) holds with $L_{f}=3c^{2}(p^{*}-1)$ and $\gamma=3$. In combination with (5.4), $L_{f}\in(0,\lambda)$ and $p^{*}>5\gamma$, we choose
@@
One can check that the associated period is $1$ and Assumptions 2.1 and 4.2 are fulfilled. It is worth noting that (5.5) has been considered in [22] to test the periodicity and mean square convergence of the backward Euler method. Thus for simplicity, we only verify the mean square order of ST methods on the time interval $[-10,10]$ with $\theta=0.75,1$ under the above settings for convergence test of (5.1). Figure 5 shows that the slopes of the error lines and the reference lines match well, indicating that the ST methods have strong rates of order one in additive case. Actually, least square fits produce rate $1.0289$ for $\theta=0.75$ and rate $1.0289$ for $\theta=1$. These facts coincide with the previous theoretical result in Theorem 4.5.